A music video website received 5,000 comments on a new song they released. The next day, the artist performed the song on television and an additional 1,500 comments were made on the website. What was the percent of increase? Round the percent to the nearest tenth if necessary.
1
Expert's answer
2020-10-06T18:04:49-0400
"\dfrac{1500}{5000}=\dfrac{x\%}{100\%}"

"x=\dfrac{1500}{5000}\cdot100\%"

"x=30\%"

The percent of increase was 30%.